# Empirical Bayes Estimation in Heterogeneous Coefficient Panel Models

### v1

@@ 3 Identification
Assumption 2(i) requires $\boldsymbol{X}$} to have full column rank, paralleling the familiar rank condition in regression. Identifying $\beta_{i}$ is generally infeasible unless some ${\boldsymbol{X}}\in\operatorname{supp}(X_{i})$ has full column rank. This ensures the existence of an annihilator matrix $\boldsymbol{M}$ with rank $T-{d_{\beta}}$. [^4] The choice of $\boldsymbol{M}$ is unique (up to a nonsingular linear transformation) given $\boldsymbol{X}$}, since a different choice of $\boldsymbol{M}$ can be written as $\tilde{\boldsymbol{M}}=\boldsymbol{M}A$ for some nonsingular $A$. We further require that ${d_{\beta}}\leq T-1$ so that $\boldsymbol{M}$ has non-null rank and can be used in part (ii). This restriction is needed because, in our setting, the covariances $\Sigma_{i}$ depend on unknown $\delta_{i}$.
Assumption 2(ii) imposes that $\delta\mapsto\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}$ is a one-to-one function of $\delta\in\mathcal{K}_{\delta}$, meaning that the variance parameter remains identifiable after removing the effect of ${\boldsymbol{X}}\beta_{i}$. The intuition is as follows. Since the concatenated matrix $({\boldsymbol{X}},\boldsymbol{M})$ is of full rank, the information in $Y_{i}$, conditional on $X_{i}={\boldsymbol{X}}$, can be expressed as
@@
where $\boldsymbol{M}^{\prime}Y_{i}$ is a Gaussian mixture with mean zero and variance $\boldsymbol{M}^{\prime}P_{i}P_{i}^{\prime}\boldsymbol{M}=\boldsymbol{M}^{\prime}P(\delta_{i})P(\delta_{i})^{\prime}\boldsymbol{M}$. The distribution of $\boldsymbol{M}^{\prime}P(\delta_{i})P(\delta_{i})^{\prime}\boldsymbol{M}$ can therefore be identified from the distribution of $\boldsymbol{M}^{\prime}Y_{i}$. Assumption 2(ii) states that $\delta_{i}$ is identified since $\delta\mapsto\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}$ is one-to-one.
@@
as a perturbation of $\boldsymbol{M}^{\prime}Y_{i}$ with noise $\mathcal{E}^{\prime}{\boldsymbol{X}}^{\prime}Y_{i}$ injected. We measure the shift in the mixing distribution of $\boldsymbol{M}^{\prime}Y_{i}+\mathcal{E}^{\prime}{\boldsymbol{X}}^{\prime}Y_{i}$ relative to the mixing distribution of $\boldsymbol{M}^{\prime}Y_{i}$ identified earlier. This comparison reveals the conditional distribution of $\mathcal{E}^{\prime}({\boldsymbol{X}}^{\prime}{\boldsymbol{X}})\beta_{i}$ given $\delta_{i}$ as $\mathcal{E}$ traverses $\mathbb{R}^{T\times(T-{d_{\beta}})}$, which then identifies $\beta_{i}$ given $\delta_{i}$ via the Cramer-Wold device.
The following theorem formalizes this result. The proof of Theorem 1 as well as the proofs of all other theorems and propositions are given in Appendix A.
@@
The primary challenge in identifying $G_{*}$ is to distinguish the mean and variance components that are convolved into the outcome distribution. For instance, assume that $Y_{i}\sim\mathcal{N}(0,1)$ is a scalar observation. The observed distribution of $Y_{i}$ can be rationalized by a continuum of models: $Y_{i}=a_{i}+\sigma_{i}e_{i}$, where $a_{i}\sim\mathcal{N}(0,v_{a})$ and $\sigma_{i}\equiv\sqrt{1-v_{a}}$ for $v_{a}\in[0,1]$. Thus, one cannot tell which model generates the observed distribution without imposing further restrictions on $a_{i}$ or $\sigma_{i}$. Bruni and Koch (1985) establish general identification results for Gaussian mixtures under a compact support assumption, which requires both the mean and variance components to be compactly supported. [^5] These results have recently attracted attention in the identification of latent distributions in learning models within labor economics (see, e.g., Pastorino (2024); Bunting, Diegert, and Maurel (2024); de Paula, Gualdani, Pastorino, and Salgado (2025)). This branch of the literature does not overlap with the broader econometrics literature on the identification of mixture models (see Compiani and Kitamura (2016) for a review). Using a type of identification-at-infinity argument, they show that both components can be separated by leveraging variation in the tails of $Y_{i}$. However, this argument breaks down as soon as the mean component places a slight nonzero mass in the tails, producing observationally indistinguishable yet fundamentally different models, as illustrated above. In contrast, our identification strategy rests on the rank and one-to-one restrictions that preclude this type of identification failure.
@@ 3.1.1 AR(1) Errors
Recall the form of the covariance matrix $\Sigma_{i}$ given in (3). This yields the likelihood function evaluated at a generic parameter $\theta=(a,\sigma^{2},\rho)$:
$\ell(Y_{i}\mid\theta)=\frac{(1-\rho^{2})^{1/2}}{(2\pi\sigma^{2})^{T/2}}\exp\left(-\frac{(1-\rho^{2})(Y_{i1}-a)^{2}}{2\sigma^{2}}-\sum_{t=2}^{T}\frac{(Y_{it}-\rho Y_{it-1}-(1-\rho)a)^{2}}{2\sigma^{2}}\right).$ (13)
We establish the identification of $G_{*}$ in the following proposition.
@@
Assume that $T\geq 3$. Then, $G_{*}$ is identified in $\mathcal{G}=\mathcal{P}(\Theta)$, where $\Theta=\mathbb{R}\times\mathcal{K}_{(\sigma^{2},\rho)}$ and $\mathcal{K}_{(\sigma^{2},\rho)}$ is a compact subset of $(0,\infty)\times(-1,1)$.
The restriction on $\mathcal{K}_{(\sigma^{2},\rho)}$ implies that $\rho_{i}$ is bounded away from $\{-1,1\}$ and that $\sigma_{i}^{2}$ is bounded away from 0 and infinity. The condition $T\geq 3$ is not only sufficient but also necessary for the identification of the three-dimensional parameter $\theta_{i}$. The variance parameters are identified from the covariance structure after purging the fixed effect $a_{i}$ via first-differencing $Y_{i}$. Finally, the compact support for $\delta_{i}$ ensures that $Y_{i}$ has non-degenerate and bounded covariance matrices across all $\delta_{i}$; in particular, this condition rules out unit root processes.
We conclude this subsection by illustrating the choice of $({\boldsymbol{X}},\boldsymbol{M})$ for Assumption 2 in this example. For simplicity, consider $T=3$. Let $\boldsymbol{X}$} be the column vector of ones whose rank is $1$. Set $\boldsymbol{M}$ equal to the first-difference matrix:
@@ 3.1.2 ARMA(1,1) Errors
Let $G_{*}(\rho_{i}+\varphi_{i}=0)=0$ denote that the probability that $\rho_{i}+\varphi_{i}=0$ equals zero under $G_{*}$.
@@
Let $\theta_{i}=(a_{i},\sigma_{i}^{2},\rho_{i},\varphi_{i})$ collect all heterogeneous coefficients with $\theta_{i}\overset{\mathrm{i.i.d.}}{\sim}G_{*}$. Assume that $T\geq 4$ and $G_{*}(\rho_{i}+\varphi_{i}=0)=0$. Then $G_{*}$ is identified in $\mathcal{G}=\mathcal{P}(\Theta)$, where $\Theta=\mathbb{R}\times\mathcal{K}_{(\sigma^{2},\rho,\varphi)}$ and $\mathcal{K}_{(\sigma^{2},\rho,\varphi)}$ is a compact subset of $(0,\infty)\times(-1,1)\times[-1,1]$.
The condition $T\geq 4$ is necessary to observe the process over a horizon sufficient to distinguish the two sources of persistence, namely $\rho_{i}$ and $\varphi_{i}$. The restriction $G_{*}(\rho_{i}+\varphi_{i}=0)=0$ ensures that the two dynamic components do not perfectly offset one another. We provide the specific choice of $({\boldsymbol{X}},\boldsymbol{M})$ for Assumption 2 in Appendix B.1.
@@ 3.2 Identification of the HIVDX Model
We now return to the HIVDX model given in (1). In this model, the parameters are partitioned as $\beta_{i}=(a_{i},b_{i})$ and $\delta_{i}=(\sigma_{i}^{2},\rho_{i})$, yielding the full parameter vector $\theta_{i}=(a_{i},b_{i},\sigma_{i}^{2},\rho_{i})$ with dimension ${d_{\theta}}=4$. Define $\Delta x_{t}:=x_{t}-x_{t-1}$ as the first difference operator applied to the sequence $(x_{t})_{t=1}^{T}$. Whether and how $G_{*}$ can be identified depends critically on the specific time-series variation of the covariates. We provide two separate identification results: one for $T=4$ and another for $T\geq 5$, which require distinct assumptions on the covariate $X_{2,it}$.
@@ 3.2.1 Case I
Suppose that $X_{2,it}$ satisfies either $\Delta X_{2,i2}=\pm\Delta X_{2,i3}\neq 0$ or $\Delta X_{2,i2}\neq\Delta X_{2,i4}$ with positive probability. For example, $X_{2,it}=X_{2,i1}+(t-1)$ is a linear trend with a random initial value. Then, we have $\Delta X_{2,it}\equiv 1$ for all $t$.
@@
Assume that $T=4$ and either $\Delta X_{2,i2}=\pm\Delta X_{2,i3}\neq 0\text{ or }\Delta X_{2,i2}\neq\Delta X_{2,i4}$ holds with positive probability. Then, $G_{*}$ is identified in $\mathcal{G}=\mathcal{P}(\Theta)$, where $\Theta=\mathbb{R}\times\mathcal{K}_{(\sigma^{2},\rho)}$ and $\mathcal{K}_{(\sigma^{2},\rho)}$ is a compact subset of $(0,\infty)\times(-1,1)$.
To illustrate, consider the case when $\Delta X_{2,it}\equiv 1$. We need to find an appropriate $({\boldsymbol{X}},\boldsymbol{M})$ Pick $X_{2,1}$ in $\operatorname{supp}(X_{2,i1})$ and define
@@
This shows that the mapping $\delta\mapsto\mathcal{V}$ is one-to-one, thereby establishing Assumption 2(ii).
@@ 3.2.2 Case II
Suppose now that $\Delta X_{2,it}$ is not identically zero for some $t\in\{2,\ldots,T\}$. For instance, consider the case where $X_{2,it}$ follows a one-time universal level shift: $X_{2,it}=\mathbf{1}\{t\geq 3\}$, such as a nationwide policy change. Appendix B.1 shows why this pattern fails to satisfy Proposition 3. However, the model is identified, provided $T\geq 5$, as we now show.
@@
To illustrate, suppose that $X_{2,it}=\mathbf{1}\{t\geq 3\}$. Then, the sequence of first differences is
$(\Delta X_{2,i2},\Delta X_{2,i3},\Delta X_{2,i4},\Delta X_{2,i5})=(0,1,0,0).$
Then, $\boldsymbol{M}$ can now be constructed as
@@
we can identify $\rho$ via the ratio $2\mathcal{V}_{23}/\mathcal{V}_{11}+1$, and subsequently recover $\sigma^{2}$. This establishes that the mapping $\delta\mapsto\mathcal{V}$ is one-to-one, as required by Assumption 2(ii). By Proposition 4, this result extends to the more general specification
$Y_{it}=a_{i}+b_{i}\,\mathbf{1}\{t\geq s\}+u_{it},\quad t=1,\ldots,5,$
for any known switch point $s\geq 2$, provided the data cover both pre- and post-treatment periods.
As two examples illustrate, incorporating covariates necessitates a significantly more involved analysis. This is because the transformation matrix $\boldsymbol{M}$ forces reliance on differenced data, which often obscures the specific autocovariance structure required to identify $\delta_{i}$. It is worth emphasizing that while the sufficient conditions in Proposition 4 are simple to verify, the underlying proof is combinatorially complex. Establishing this general result requires an exhaustive analysis of all distinct patterns of $\Delta X_{2,it}$ to verify the one-to-one mapping between $\delta$ and $\mathcal{V}$ in every possible scenario.
@@ 4 Consistency of the NPMLE
Alternative metrics are discussed in the NPMLE literature, though they serve theoretical roles distinct from our metric $d$. The first is the Hellinger distance between the marginal densities, $\frac{1}{2}\int(\sqrt{f_{\hat{G}}(y)}-\sqrt{f_{G_{*}}(y)})^{2}dy$. While Hellinger distance is frequently considered in mixture models (Jiang and Zhang, 2009; Soloff, Guntuboyina, and Sen, 2025), it measures proximity in the data space rather than the parameter space. In our setting, the convergence of the fitted marginal density $f_{\hat{G}}$ to the truth $f_{G_{*}}$ is of secondary importance. Another alternative is the Wasserstein distance (e.g., $W_{2}$ as in Soloff, Guntuboyina, and Sen, 2025). However, $W_{2}$ requires the convergence of second moments and consequently becomes unbounded if probability mass escapes to infinity. This renders it unsuitable for our initial goal of establishing consistency through compactification. By contrast, our metric $d$ allows us to successfully apply the Kiefer-Wolfowitz machinery. It strikes a necessary balance: it induces a topology sufficiently weak to ensure compactness (accommodating mass at infinity), yet strong enough to imply the consistency of downstream EB estimation.

### v2

@@ 3 Identification
Assumption 2(i) requires ${\boldsymbol{X}}\in\operatorname{supp}(X_{i})$ to have full column rank, paralleling the familiar rank condition in regression. This ensures the existence of an annihilator matrix $\boldsymbol{M}$ with rank $T-{d_{\beta}}$. [^4] The choice of $\boldsymbol{M}$ is unique (up to a nonsingular linear transformation) given $\boldsymbol{X}$}, since a different choice of $\boldsymbol{M}$ can be written as $\tilde{\boldsymbol{M}}=\boldsymbol{M}A$ for some nonsingular $A$. We further require that ${d_{\beta}}\leq T-1$ so that $\boldsymbol{M}$ has non-null rank and can be used in part (ii). This restriction is needed because, in our setting, the covariances $\Sigma_{i}$ depend on unknown $\delta_{i}$.
Assumption 2(ii) imposes that $\delta\mapsto\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M}$ is a one-to-one function of $\delta\in\mathcal{K}_{\delta}$, meaning that the variance parameters remain identifiable after removing the effect of ${\boldsymbol{X}}\beta_{i}$. The intuition is as follows. Since the concatenated matrix $({\boldsymbol{X}},\boldsymbol{M})$ is of full rank, the information in $Y_{i}$, conditional on $X_{i}={\boldsymbol{X}}$, can be expressed as
@@
where $\boldsymbol{M}^{\prime}Y_{i}$ is a Gaussian mixture with mean zero and variance $\boldsymbol{M}^{\prime}P_{i}P_{i}^{\prime}\boldsymbol{M}=\boldsymbol{M}^{\prime}P(\delta_{i})P(\delta_{i})^{\prime}\boldsymbol{M}$. The distribution of $\boldsymbol{M}^{\prime}P(\delta_{i})P(\delta_{i})^{\prime}\boldsymbol{M}$ can therefore be identified from the distribution of $\boldsymbol{M}^{\prime}Y_{i}$. Then, the distribution of $\delta_{i}$ is identified in view of Assumption 2(ii).
@@
as a perturbation of $\boldsymbol{M}^{\prime}Y_{i}$ with noise $\mathcal{E}^{\prime}{\boldsymbol{X}}^{\prime}Y_{i}$. We measure the shift in the mixing distribution of $\boldsymbol{M}^{\prime}Y_{i}+\mathcal{E}^{\prime}{\boldsymbol{X}}^{\prime}Y_{i}$ relative to the mixing distribution of $\boldsymbol{M}^{\prime}Y_{i}$ identified earlier. This comparison reveals the conditional distribution of $\mathcal{E}^{\prime}({\boldsymbol{X}}^{\prime}{\boldsymbol{X}})\beta_{i}$ given $\delta_{i}$ as $\mathcal{E}$ traverses $\mathbb{R}^{T\times(T-{d_{\beta}})}$, which then identifies $\beta_{i}$ given $\delta_{i}$ via the Cramer-Wold device.
The following theorem formalizes this result. The proof of Theorem 1 as well as the proofs of all other theorems and propositions are given in Appendix S-1.
@@
The primary challenge in identifying $G_{*}$ is to distinguish the mean and variance components that are convolved into the outcome distribution. For instance, assume that $Y_{i}\sim\mathcal{N}(0,1)$ is a scalar observation. The observed distribution of $Y_{i}$ can be rationalized by a continuum of models: $Y_{i}=a_{i}+\sigma_{i}e_{i}$, where $a_{i}\sim\mathcal{N}(0,v_{a})$ and $\sigma_{i}\equiv\sqrt{1-v_{a}}$ for $v_{a}\in[0,1]$. Thus, one cannot tell which model generates the observed distribution without imposing further restrictions on $a_{i}$ or $\sigma_{i}$. Bruni and Koch (1985) establish general identification results for Gaussian mixtures under a compact support assumption for both the mean and variance components. [^5] The results of Bruni and Koch (1985) are also used in identification of latent distributions in labor economic analyses (see, e.g., Pastorino (2024); Bunting, Diegert, and Maurel (2024); de Paula, Gualdani, Pastorino, and Salgado (2025)), but the context is different from identification of mixture models in econometrics such as discussed in Compiani and Kitamura (2016). Using a type of identification-at-infinity argument, they show that both components can be separated by leveraging variation in the tails of $Y_{i}$. However, this argument breaks down as soon as the mean component places a slight nonzero mass in the tails, producing observationally indistinguishable yet fundamentally different models, as illustrated above. In contrast, our identification strategy rests on the rank and one-to-one restrictions that preclude this type of identification failure.
@@ 3.1.1 AR(1) Errors
Assume that $T\geq 3$. Consider the panel AR(1) model given in 12. $G_{*}$ is identified in $\mathcal{G}=\mathcal{P}(\Theta)$, where $\Theta=\mathbb{R}\times\mathcal{K}_{(\sigma^{2},\rho)}$ and $\mathcal{K}_{(\sigma^{2},\rho)}$ is a compact subset of $(0,\infty)\times(-1,1)$.
The restriction on $\mathcal{K}_{(\sigma^{2},\rho)}$ implies that $\rho_{i}$ is bounded away from $\{-1,1\}$ and that $\sigma_{i}^{2}$ is bounded away from 0 and infinity. The condition $T\geq 3$ is not only sufficient but also necessary for the identification of the three-dimensional parameter $\theta_{i}$. The variance parameters are identified from the covariance structure after purging the fixed effect $a_{i}$ by first-differencing $Y_{i}$ via a suitably chosen $\boldsymbol{M}$. Finally, the compact support for $\delta_{i}$ ensures that $Y_{i}$ has non-degenerate and bounded covariance matrices across all $\delta_{i}$; in particular, this condition rules out unit root processes. Our choice of $\boldsymbol{M}$ can be illustrated with $T=3$. Let $\boldsymbol{X}$} be the column vector of ones whose rank is $1$. Set $\boldsymbol{M}$ equal to the first-difference matrix:
@@ 3.1.2 ARMA(1,1) Errors
Let $\theta_{i}=(a_{i},\sigma_{i}^{2},\rho_{i},\varphi_{i})$ collect all heterogeneous coefficients in the panel model with ARMA(1,1) errors, with $\theta_{i}\overset{\mathrm{i.i.d.}}{\sim}G_{*}$. Assume that $T\geq 4$ and let $G_{*}(\rho_{i}+\varphi_{i}=0)=0$ denote that the probability that $\rho_{i}+\varphi_{i}=0$ equals zero under $G_{*}$. Then $G_{*}$ is identified in $\mathcal{G}=\mathcal{P}(\Theta)$, where $\Theta=\mathbb{R}\times\mathcal{K}_{(\sigma^{2},\rho,\varphi)}$ and $\mathcal{K}_{(\sigma^{2},\rho,\varphi)}$ is a compact subset of $(0,\infty)\times(-1,1)\times[-1,1]$.
The condition $T\geq 4$ is necessary to observe the process over a horizon sufficient to distinguish the two sources of persistence, namely $\rho_{i}$ and $\varphi_{i}$. The restriction $G_{*}(\rho_{i}+\varphi_{i}=0)=0$ ensures that the two dynamic components do not perfectly offset one another. We provide the specific choice of $({\boldsymbol{X}},\boldsymbol{M})$ for Assumption 2 in Appendix O-1.
@@ 3.2 Identification of the HIVDX Model
We now return to the HIVDX model given in (1). In this model, the parameters are partitioned as $\beta_{i}=(a_{i},b_{i})$ and $\delta_{i}=(\sigma_{i}^{2},\rho_{i})$, yielding the full parameter vector $\theta_{i}=(a_{i},b_{i},\sigma_{i}^{2},\rho_{i})$ with dimension ${d_{\theta}}=4$. Define $\Delta x_{t}:=x_{t}-x_{t-1}$ as the first difference operator applied to the sequence $(x_{t})_{t=1}^{T}$. Whether and how $G_{*}$ can be identified depends critically on the specific time-series variation of the covariates. We illustrate this using two processes for $X_{2it}$. In the first case, the parameters of the HIVDX model are identified when $T\geq 4$, whereas identification requires $T\geq 5$ in the second case.
@@ 3.2.1 Case I
We start with the following proposition.
@@
Assume that $T\geq 4$ and either $\Delta X_{2,i2}=\pm\Delta X_{2,i3}\neq 0\text{ or }\Delta X_{2,i2}\neq\Delta X_{2,i4}$ holds with positive probability. Then, $G_{*}$ is identified in $\mathcal{G}=\mathcal{P}(\Theta)$, where $\Theta=\mathbb{R}\times\mathcal{K}_{(\sigma^{2},\rho)}$ and $\mathcal{K}_{(\sigma^{2},\rho)}$ is a compact subset of $(0,\infty)\times(-1,1)$.
To illustrate, suppose that $X_{2,it}=X_{2,i1}+(t-1)$ is a linear trend with a random initial value. Then, we have $\Delta X_{2,it}\equiv 1$ for all $t$. We need to find an appropriate $({\boldsymbol{X}},\boldsymbol{M})$. Pick $X_{2,1}$ in $\operatorname{supp}(X_{2,i1})$ and define
@@
This shows that the mapping $\delta\mapsto\mathcal{V}$ is one-to-one , thereby satisfying Assumption 2(ii). However, as shown in Appendix O-1, Proposition 3 will not identify the $X_{2}$ process in Case II that will now be discussed.
@@ 3.2.2 Case II
An example of $X_{2,it}$ that satisfies the condition in the proposition is a one-time universal level shift: $X_{2,it}=\mathbf{1}\{t\geq 3\}$, as in a nationwide policy change. Then, the sequence of first differences is $(\Delta X_{2,i2},\Delta X_{2,i3},\Delta X_{2,i4},\Delta X_{2,i5})=(0,1,0,0).$ Then, $\boldsymbol{M}$ can now be constructed as
@@
we can identify $\rho$ via the ratio $2\mathcal{V}_{23}/\mathcal{V}_{11}+1$, and subsequently recover $\sigma^{2}$. This establishes that the mapping $\delta\mapsto\mathcal{V}$ is one-to-one, as required by Assumption 2(ii). The result in Proposition 4 also extends to the analysis of heterogeneous treatment implemented at a known time $s\geq 2$, giving
$Y_{it}=a_{i}+b_{i}\,\mathbf{1}\{t\geq s\}+u_{it},\quad t=1,\ldots,5.$
Then the term $\mathbf{1}\{t\geq s\}$ represents a common treatment or policy, and $b_{i}$ captures the individual-specific treatment effect.
As these examples illustrate, incorporating covariates necessitates a significantly more involved analysis. This is because the transformation matrix $\boldsymbol{M}$ that yields differenced data often obscures the autocovariance structure required to identify $\delta_{i}$. It is worth emphasizing that while the sufficient conditions in Proposition 4 are simple to verify, the underlying proof is combinatorially complex. Establishing this general result requires an exhaustive analysis of all distinct patterns of $\Delta X_{2,it}$ to verify the one-to-one mapping between $\delta$ and $\mathcal{V}$ in every possible scenario, as shown in the Appendix.
@@ 4 Consistency of the NPMLE
Alternative metrics are discussed in the NPMLE literature, though they serve theoretical roles distinct from our metric $d$. The first is the Hellinger distance between the marginal densities, $\frac{1}{2}\int(\sqrt{f_{\hat{G}}(y)}-\sqrt{f_{G_{*}}(y)})^{2}dy$. While Hellinger distance is frequently considered in mixture models (Jiang and Zhang, 2009; Soloff, Guntuboyina, and Sen, 2025), it measures proximity in the data space rather than the parameter space. In our setting, the convergence of the fitted marginal density $f_{\hat{G}}$ to the truth $f_{G_{*}}$ is of secondary importance. Another alternative is the Wasserstein distance (e.g., $W_{2}$ as in Soloff, Guntuboyina, and Sen, 2025). However, $W_{2}$ requires the convergence of second moments and consequently becomes unbounded if probability mass escapes to infinity. This renders it unsuitable for our goal of establishing consistency through compactification. The metric $d$ allows us to apply the Kiefer-Wolfowitz machinery. It strikes a necessary balance: it induces a topology sufficiently weak to ensure compactness (accommodating mass at infinity), yet strong enough to imply the consistency of downstream EB estimation.